jon is twice as old as his sister, jane, less 4 years. jon is 18 years old. how old is jane?
Answer by cathieb1225(41)

x is Jane's age;2x-4 is Jon's age. Because he is 18 years old, the equation would be 2x-4=18.
2x-4=18
Add 4 to both sides of the equation
2x-4+4=18+4
2x=22
Divide both sides by 2
2x/2=22/2
x=11
Jane is 11 years old.
To check your answer, substitute 11 for x. 2(11)-4=18
22-4=18
18=18